A is 12 years older than B; 12 years ago he was twice as old as B then was. How old is A now, and how many years ago is it since he was three times as old as B then was

Answer by ankor@dixie-net.com(22740)

A is 12 years older than B;
a = b + 12
a - b = 12
:
12 years ago he was twice as old as B then was.
a - 12 = 2(b - 12)
a - 12 = 2b - 24
a - 2b = -24 + 12
a - 2b = -12
use elimination on these two equations
a - b = 12
a - 2b =-12
----------------subtraction eliminates a, find b
0 + b = 24
B is 24 yrs old
:
How old is A now,
"A is 12 years older than B"
a = 12 + 24
a = 36 yrs is A
:
how many years ago is it since he was three times as old as B then was
let y = no. of years for this to be true
36 - y = 3(24 - y)
36 - y = 72 - 3y
-t + 3y = 72 - 36
2y = 36
y = 36/2
y = 18 yrs ago A was 3 times as old a B